# GeoGen: A Two-stage Coarse-to-Fine Framework for Fine-grained Synthetic Location-based Social Network Trajectory Generation

### v1

@@ Introduction
SMM (Maglaras and Katsaros 2015) 0.196 0.259 0.381 0.358 0.178 0.277 0.217 0.329
TimeGEO (Jiang et al. 2016) 0.251 0.625 0.278 0.386 0.277 0.524 0.232 0.216
Hawkes (Laub, Taimre, and Pollett 2015) 0.190 0.517 0.339 0.253 0.157 0.315 0.269 0.281
LSTM (Rossi et al. 2021) 0.166 0.441 0.175 0.235 0.184 0.292 0.230 0.257
SeqGAN (Yu et al. 2017) 0.116 0.364 0.150 0.291 0.108 0.292 0.207 0.276
MoveSim (Feng et al. 2020) 0.080 0.289 0.133 0.391 0.079 0.256 0.114 0.353
DiffTraj (Zhu et al. 2023b) 0.051 0.085 0.104 0.320 0.276 0.201 0.116 0.744
ControlTraj (Zhu et al. 2024) 0.193 0.356 0.160 0.337 0.049 0.096 0.191 0.223
Our Method (ours) 0.079 0.065 0.067 0.127 0.015 0.043 0.081 0.120
@@
SMM (Maglaras and Katsaros 2015) 0.603 0.625 0.139 0.347 0.637 0.599 0.697 0.428
TimeGEO (Jiang et al. 2016) 0.416 0.501 0.148 0.401 0.579 0.489 0.727 0.397
Hawkes (Laub, Taimre, and Pollett 2015) 0.517 0.556 0.261 0.521 0.621 0.496 0.579 0.433
LSTM (Rossi et al. 2021) 0.589 0.602 0.187 0.339 0.618 0.527 0.699 0.369
SeqGAN (Yu et al. 2017) 0.452 0.475 0.029 0.292 0.531 0.462 0.745 0.413
MoveSim (Feng et al. 2020) 0.306 0.382 0.078 0.275 0.181 0.265 0.318 0.138
DiffTraj (Zhu et al. 2023b) 0.061 0.104 0.241 0.702 0.114 0.275 0.802 0.306
ControlTraj (Zhu et al. 2024) 0.108 0.217 0.150 0.286 0.204 0.253 0.765 0.141
Our Method (ours) 0.058 0.088 0.036 0.200 0.136 0.197 0.292 0.080
@@
The Multi-task Hybrid-Head Decoder autoregressively generates the fine-grained trajectory, i.e., a sequence of (POI, timestamp) pairs, conditioned on the contextual representations $H$ from the encoder. To generate the $n^{\prime}$-th check-in point, the previously generated sequence $s^{\prime}_{\mathit{syn}}=[(p_{1},t_{1}),\ldots,(p_{n^{\prime}-1},t_{n^{\prime}-1})]$ is first embedded into a sequence of vectors $s^{\prime}_{vec}$, where each POI is replaced by its shared embedding and each timestamp is converted using Time2Vec. Next, a Transformer-based Decoder processes this embedded sequence and the encoder’s output $H$ to produce a hidden state $h_{n^{\prime}}$, which summarizes the necessary history and context. This hidden state is then fed into two specialized prediction heads to determine where and when the next check-in occurs. For POI prediction, a linear layer computes logits over the vocabulary: $\text{logits}_{\text{POI}}=W_{\text{POI}}h_{n^{\prime}}+b_{\text{POI}}$. For timestamp prediction, we model the temporally irregular nature of check-ins using a Neural Temporal Point Process (NTPP)  (Shchur et al. 2021). Unlike direct regression, which provides a single, often unrealistic point estimate by simply predicting an average time interval. The TPP models the full probability distribution of the next event by predicting the conditional intensity function $\lambda(t)=\text{Softplus}(W_{\text{time}}h_{n^{\prime}}+b_{\text{time}})$, which defines the instantaneous event rate given the current context. To sample the next timestamp, we first derive the cumulative distribution function $F^{-1}$ corresponding to $\lambda(t)$, and then draw a sample from this distribution via inverse transform sampling:
@@
The POI prediction loss $\mathcal{L}_{\text{POI}}=-\sum_{t=1}^{T}y_{t}\log(\hat{y}_{t})$ enforces categorical accuracy through cross-entropy, where $T$ represents the number of check-ins in a trajectory and $y_{t}$, $\hat{y}_{t}$ denote ground-truth and predicted POI id respectively. The temporal loss is the negative log-likelihood objective for the neural temporal point processes, formulated as $\mathcal{L}_{\text{time}}=-\sum_{i=1}^{T}(\log\lambda_{i}-\lambda_{i}\Delta t_{i})$ , where $\lambda_{i}$ is the intensity predicted by the network for the $i$-th interval, and $\Delta t_{i}$ is the ground-truth time duration since the last event. This loss maximizes the probability of the observed event sequence by training the model to predict a high intensity just before an event occurs. The spatial consistency loss $\mathcal{L}_{\text{spatial}}=\frac{1}{T}\sum_{t=1}^{T}\|p_{t}-p_{\hat{t}}\|_{2}$ penalizes geographic implausibility by measuring Euclidean distance between predicted and ground-truth POI locations, where $p_{t}$ and $p_{\hat{t}}$ denote the geographic coordinates of the ground-truth and predicted POIs, respectively. The three components are balanced by three learnable weights, enabling the model to generate trajectories that are spatially plausible, temporally accurate, and categorically consistent.
@@
Figure 4: Performance comparison of different methods on FS-NYC dataset.
@@
We utilize four publicly available LBSN datasets containing POI check-in trajectories to evaluate the performance of our proposed GeoGen, including Foursquare New York City (FS-NYC), Foursquare Tokyo (FS-TKY), Foursquare Austin (FS-ATX) (Yang et al. 2014; Yang, Zhang, and Qu 2016), and Gowalla (GW-STO) (Cho, Myers, and Leskovec 2011). We first truncate each individual LBSN check-in trajectory with a pre-defined duration (one week for the three FourSquare datasets and six weeks for GW-STO) and maintain the sequences with more than 10 visits to ensure data quality. Detailed data descriptions are shown in the Appendix. We split each dataset into training, validation, and testing sets with a ratio of 7:2:1.
@@
We compare our GeoGen with the following eight state-of-the-art models: SMM (Maglaras and Katsaros 2015), TimeGEO (Jiang et al. 2016), Hawkes (Laub, Taimre, and Pollett 2015), LSTM (Rossi et al. 2021), SeqGAN (Yu et al. 2017), MoveSim (Feng et al. 2020), DiffTraj (Zhu et al. 2023b), ControlTraj (Zhu et al. 2024). Detailed descriptions of these baselines are shown in the Appendix.
@@
We evaluate the synthetic data quality from both fidelity and utility perspectives. Following common practice (Zhu et al. 2023b; Feng et al. 2020), we evaluate data fidelity by computing the Jensen-Shannon Divergence (JSD) (Lin 1991) between the distributions of synthetic and real data across four metrics from both spatial and temporal aspects: 1) Distance: the distance between consecutive check-ins; 2) Radius: the radius of gyration, which measures the trajectory’s spatial range; 3) Interval: the time interval between consecutive check-ins; 4) Length: the total number of points in the trajectory. A lower JSD score indicates a higher fidelity of the synthetic data. We evaluate the utility of synthetic data by conducting data-driven downstream applications.
@@
The comprehensive quantitative results are presented in Table 1. We found that traditional rule-based methods (SMM, TimeGEO, Hawkes) generally fall short, as their reliance on fixed statistical assumptions prevents them from capturing the complex, long-term behaviors from POI check-in trajectories. Designed for spatially continuous and temporally regular trajectory data generation, DiffTraj shows competitive performance on spatial metrics on dense urban datasets like FS-NYC, but performs poorly on the length metric due to the lack of awareness of discrete POI locations. Although the ControlTraj can model the connection between POIs as road segments, its incompatibility with temporally irregular data degrades the LBSN check-in trajectory generation quality. In contrast, our proposed model, GeoGen, significantly outperforms baselines across the four diverse datasets, highlighting its robustness and generalizability to generate high-fidelity LBSN trajectory data. For instance, on the FS-TKY dataset, GeoGen improves the second-best method by over 69% (0.049 - 0.015 / 0.049) in the Distance metric and 55% in the Radius metric. Similarly, it achieves a 38% improvement in trajectory Length on the GW-STO dataset.
The Cumulative Distribution Function (CDF) curves in Figure 4 (and more results in the Appendix) show that GeoGen’s distributions most closely align with those of the original data, confirming its effectiveness in capturing realistic check-in behavior patterns.
Table 2: Synthetic data utility for next check-in prediction.
@@
Utility Evaluation
As shown in Table 2, we evaluate the utility of the generated synthetic data through a next check-in prediction task (Yuan et al. 2023), where a prediction model is trained on synthetic data and tested on real-world data. Root Mean Square Error (RMSE) and Euclidean Distance (ED) are used to measure the prediction performance from the temporal dimension and spatial dimension, respectively. Our results demonstrate that GeoGen achieves the closest performance to the real data across both temporal and spatial metrics using FS-NYC and FS-TKY datasets, indicating its strong ability to preserve data utility. MoveSim performs better than SeqGAN and ControlTraj in the spatial dimension due to its explicit consideration of spatial structures such as distances between locations, while SeqGAN and ControlTraj treat locations as discrete categorical variables. Our proposed GeoGen consistently achieves the best performance across all metrics and datasets, thanks to its ability to effectively model spatially discrete and temporally irregular patterns of LBSN check-in trajectories.
@@
We investigate the impact of time intervals $I$. As shown in Table 4, we found that smaller intervals (e.g., 2-4 hours) yield better temporal fidelity (Interval), whereas spatial fidelity metrics (Distance and Radius) are optimized at a larger 6-hour interval. Increasing the interval from 2 to 6 hours reduces Stage 1 memory consumption from $2.36$ GB to $0.35$ GB and increases throughput by more than fourfold (from $11,322$ to $50,540$ samples/sec). In the second stage, the computational cost of Coarse2FineNet remains relatively stable across different intervals.
Table 4: Results under different coarse-grained interval II selection on GW-STO dataset with batch size 256 for S2TDiff (S1) and 16 for Coarse2FineNet (S2). The unit of throughput is samples per second.
@@
The recent advances in generative models provide a great opportunity for synthetic trajectory data generation. These models (e.g., Generative Adversarial Network (GAN) (Rao et al. 2020; Liu, Chen, and Andris 2018; Cao and Li 2021) and diffusion models (Zhu et al. 2023b)) utilize the power of deep neural networks to learn and replicate the underlying spatio-temporal distributions of real trajectory data. Once trained, these models generate synthetic trajectories by sampling from the learned distributions. For example, TrajGen (Cao and Li 2021) utilizes GAN and a Seq2Seq model to generate spatial-temporal trajectory data. Feng et al. (Feng et al. 2020) propose a GAN-based framework, MoveSim, with the introduction of prior knowledge of human mobility. MoveSim (Feng et al. 2020) generates the next location based on the last point in the partially generated sequence and iteratively generates the rest trajectory points. Zhu et al. (Zhu et al. 2023b, a, 2024) propose a pioneering work by adopting the diffusion model to generate continuous and road network-constrained trajectories with a fixed length and time interval between two trajectory points. Different from existing works that usually focus on generating trajectories in continuous spatial spaces with fixed time intervals and fixed lengths, our work aims to generate spatially discrete, temporally irregular LBSN check-in trajectories.

### v2

@@ Introduction
SMM (Maglaras and Katsaros 2015) 0.196 0.259 0.381 0.358 0.299 0.178 0.277 0.217 0.329 0.250
TimeGEO (Jiang et al. 2016) 0.251 0.625 0.278 0.386 0.385 0.277 0.524 0.232 0.216 0.312
Hawkes (Laub, Taimre, and Pollett 2015) 0.190 0.517 0.339 0.253 0.325 0.157 0.315 0.269 0.281 0.256
LSTM (Rossi et al. 2021) 0.166 0.441 0.175 0.235 0.254 0.184 0.292 0.230 0.257 0.241
SeqGAN (Yu et al. 2017) 0.116 0.364 0.150 0.291 0.230 0.108 0.292 0.207 0.276 0.221
MoveSim (Feng et al. 2020) 0.080 0.289 0.133 0.391 0.223 0.079 0.256 0.114 0.353 0.201
DiffTraj (Zhu et al. 2023b) 0.051 0.085 0.104 0.320 0.140 0.276 0.201 0.116 0.744 0.334
ControlTraj (Zhu et al. 2024) 0.193 0.356 0.160 0.337 0.262 0.049 0.096 0.191 0.223 0.140
Our Method (ours) 0.079 0.065 0.067 0.127 0.085 0.015 0.043 0.081 0.120 0.065
@@
SMM (Maglaras and Katsaros 2015) 0.603 0.625 0.139 0.347 0.429 0.637 0.599 0.697 0.428 0.590
TimeGEO (Jiang et al. 2016) 0.416 0.501 0.148 0.401 0.367 0.579 0.489 0.727 0.397 0.548
Hawkes (Laub, Taimre, and Pollett 2015) 0.517 0.556 0.261 0.521 0.464 0.621 0.496 0.579 0.433 0.532
LSTM (Rossi et al. 2021) 0.589 0.602 0.187 0.339 0.429 0.618 0.527 0.699 0.369 0.553
SeqGAN (Yu et al. 2017) 0.452 0.475 0.029 0.292 0.312 0.531 0.462 0.745 0.413 0.538
MoveSim (Feng et al. 2020) 0.306 0.382 0.078 0.275 0.260 0.181 0.265 0.318 0.138 0.226
DiffTraj (Zhu et al. 2023b) 0.061 0.104 0.241 0.702 0.277 0.114 0.275 0.802 0.306 0.374
ControlTraj (Zhu et al. 2024) 0.108 0.217 0.150 0.286 0.190 0.204 0.253 0.765 0.141 0.341
Our Method (ours) 0.058 0.088 0.036 0.200 0.096 0.136 0.197 0.292 0.080 0.176
Table 1: Fidelity Evaluation. The best results on each dataset are in bold, and the second-best results are underlined. ‘Average’ denotes the mean value of the four metrics.
@@
The Multi-task Hybrid-Head Decoder autoregressively generates a sequence of (POI, timestamp) pairs, conditioned on the contextual representations $H$ from the encoder. To generate the $n^{\prime}$-th check-in point, the previously generated sequence $s^{\prime}_{\mathit{syn}}=[(p_{1},t_{1}),\ldots,(p_{n^{\prime}-1},t_{n^{\prime}-1})]$ is first embedded into a sequence of vectors $s^{\prime}_{vec}$, where each POI is replaced by its shared embedding and each timestamp is converted using Time2Vec. Next, a Transformer-based Decoder processes this embedded sequence and the encoder’s output $H$ to produce a hidden state $h_{n^{\prime}}$, which summarizes the necessary history and context. This hidden state is then fed into two specialized prediction heads to determine where and when the next check-in occurs. For POI, a linear layer computes logits over the vocabulary. For timestamps, unlike direct regression, we adopt a Neural Temporal Point Process (NTPP) (Shchur et al. 2021) to capture the irregular timing of human mobility. The NTPP models the probability distribution of the next timestamp through a conditional intensity function $\lambda(t)=\text{Softplus}(W_{\text{time}}h_{n^{\prime}}+b_{\text{time}}),$ which defines the instantaneous event rate given the current context. To sample the next timestamp, we first derive the cumulative distribution function $F^{-1}$ corresponding to $\lambda(t)$, and then draw a sample from this distribution via inverse transform sampling:
@@
The POI prediction loss $\mathcal{L}_{\text{POI}}=-\sum_{t=1}^{T}y_{t}\log(\hat{y}_{t})$ enforces categorical accuracy through cross-entropy, where $T$ represents the number of check-ins in a trajectory and $y_{t}$, $\hat{y}_{t}$ denote ground-truth and predicted POI ID respectively. The temporal loss is the negative log-likelihood objective for the neural temporal point processes, formulated as $\mathcal{L}_{\text{time}}=-\sum_{i=1}^{T}(\log\lambda_{i}-\lambda_{i}\Delta t_{i})$ , where $\lambda_{i}$ is the intensity predicted by the network for the $i$-th interval, and $\Delta t_{i}$ is the ground-truth time duration since the last check-in. This loss maximizes the probability of the observed check-in sequence by training the model to predict a high intensity just before a check-in occurs. The spatial consistency loss $\mathcal{L}_{\text{spatial}}=\frac{1}{T}\sum_{t=1}^{T}\|g_{p_{t}}-g_{p_{\hat{t}}}\|_{2}$ penalizes geographic implausibility by measuring Euclidean distance between predicted POI $p_{\hat{t}}$ and ground-truth POI $p_{t}$. The three components are balanced by three learnable weights, enabling the model to generate trajectories that are spatially plausible, temporally accurate, and categorically consistent.
@@
Figure 4: Performance comparison of different methods on FS-TKY dataset.
@@
We utilize four publicly available LBSN datasets containing POI check-in trajectories to evaluate the performance of our proposed GeoGen, including New York City (FS-NYC), Tokyo (FS-TKY), Austin (FS-ATX) from the Foursquare dataset(Yang et al. 2014; Yang, Zhang, and Qu 2016), and Stockholm (GW-STO) from the Gowalla dataset (Cho, Myers, and Leskovec 2011). We first truncate each individual LBSN check-in trajectory with a pre-defined duration (one week for the three FourSquare datasets and six weeks for GW-STO) and maintain the sequences with more than 5 visits to ensure data quality. We split each dataset into training, validation, and testing sets with a ratio of 7:1:2.
@@
We compare our GeoGen with the following eight state-of-the-art models: SMM (Maglaras and Katsaros 2015), TimeGEO (Jiang et al. 2016), Hawkes (Laub, Taimre, and Pollett 2015), LSTM (Rossi et al. 2021), SeqGAN (Yu et al. 2017), MoveSim (Feng et al. 2020), DiffTraj (Zhu et al. 2023b), ControlTraj (Zhu et al. 2024).
@@
We evaluate the synthetic data quality from both fidelity and utility perspectives. Following common practice (Zhu et al. 2023b; Feng et al. 2020), we evaluate data fidelity by computing the Jensen-Shannon Divergence (JSD) (Lin 1991) between the distributions of synthetic and real data across four metrics from both spatial and temporal aspects: 1) Distance: the sum of distance between consecutive check-ins; 2) Radius: the radius of gyration, which measures the trajectory’s spatial range; 3) Interval: the time interval between consecutive check-ins; 4) Length: the total number of points in the trajectory. A lower JSD score indicates a higher fidelity of the synthetic data. We evaluate the utility of synthetic data by conducting data-driven downstream applications.
@@
The comprehensive quantitative results are presented in Table 1. We found that traditional rule-based methods (SMM, TimeGEO, Hawkes) generally fall short, as their reliance on fixed statistical assumptions prevents them from capturing the complex, long-term behaviors from POI check-in trajectories. Designed for spatially continuous and temporally regular trajectory data generation, DiffTraj shows competitive performance on spatial metrics on dense urban datasets like FS-NYC, but performs poorly on the length metric due to the lack of awareness of discrete POI locations. Although the ControlTraj can model the connection between POIs as road segments, its incompatibility with temporally irregular data degrades the LBSN check-in trajectory generation quality. In contrast, our proposed model, GeoGen, significantly outperforms baselines across the four diverse datasets, highlighting its robustness and generalizability to generate high-fidelity LBSN trajectory data. For instance, on the FS-TKY dataset, GeoGen improves the second-best method by over 69% in the Distance metric and 55% in the Radius metric. Similarly, it achieves a 38% improvement in Length on the GW-STO dataset.
The Cumulative Distribution Function (CDF) curves in Figure 4 for FS-TKY show that GeoGen’s distributions most closely align with the original data, confirming its effectiveness in capturing realistic LBSN check-in behavior patterns.
Utility Evaluation
As shown in Table 2, we evaluate the utility of the synthetic data through a next check-in prediction task (Yuan et al. 2023), where a prediction model is trained on synthetic data and tested on real-world data. Root Mean Square Error (RMSE) and Euclidean Distance (ED) are used to measure the prediction performance from the temporal dimension and spatial dimension, respectively. Our results demonstrate that GeoGen achieves the closest performance to the real data across both temporal and spatial metrics using FS-NYC and FS-TKY datasets, indicating its strong ability to preserve data utility. MoveSim performs better than SeqGAN and ControlTraj in the spatial dimension due to its explicit consideration of spatial structures such as distances between locations, while SeqGAN and ControlTraj treat locations as discrete categorical variables. Our proposed GeoGen consistently achieves the best performance across all metrics and datasets, thanks to its ability to effectively model spatially discrete and temporally irregular patterns of LBSN check-in trajectories.
@@
Table 2: Synthetic data utility for next check-in prediction.
@@
We investigate the impact of time intervals $I$. As shown in Table 4, we found that smaller intervals (e.g., 2-4 hours) yield better temporal fidelity (Interval), whereas spatial fidelity metrics (Distance and Radius) are optimized at a larger 6-hour interval. Increasing the interval from 2 to 6 hours reduces Stage 1 memory consumption from $2.36$ GB to $0.35$ GB and increases throughput by more than fourfold (from $11,322$ to $50,540$ samples/sec). In the second stage, the computational cost of Coarse2FineNet remains relatively stable across different intervals. Consequently, a moderate $I$ strikes a balance between generation quality and efficiency.
@@
Table 4: Results under different coarse-grained interval II selection on GW-STO dataset with batch size 256 for S2TDiff (S1) and 16 for Coarse2FineNet (S2).
@@
The recent advances in generative models provide a great opportunity for synthetic trajectory data generation. These models (e.g., Generative Adversarial Network (GAN) (Rao et al. 2020; Liu, Chen, and Andris 2018; Cao and Li 2021) and diffusion models (Zhu et al. 2023b)) utilize the power of deep neural networks to learn and replicate the underlying spatio-temporal distributions of real trajectory data. Once trained, these models generate synthetic trajectories by sampling from the learned distributions. For example, Feng et al. (Feng et al. 2020) propose a GAN-based framework, MoveSim, with the introduction of prior knowledge for human mobility trajectory generation. MoveSim (Feng et al. 2020) generates the next location based on the last point in the partially generated sequence and iteratively generates the rest trajectory points. Zhu et al. (Zhu et al. 2023b, a, 2024) propose a pioneering work by adopting the diffusion model to generate continuous and road network-constrained trajectories with a fixed length and time interval between two trajectory points. Different from existing works that usually focus on generating trajectories in continuous spatial spaces with fixed time intervals and fixed lengths, our work aims to generate spatially discrete, temporally irregular LBSN check-in trajectories.